OpenStudy (anonymous):

a-2√(a)=3 how would I solve for a?

OpenStudy (anonymous):

\[\implies a-3=2\sqrt{a}\] \[(a-3)^{2}=4a\] \[a ^{2}-6a+9=4a\] \[a^{2}-10a+\] then solve for a=1 or 9